a man weighing 160 pounds stands on the floor of an elevator what force will be exerted on the floor what if it descends with a uniform acceleration of 1 ft ft per second square and ascends with a uniform acceleration of 1 ft per second square


1
Expert's answer
2022-01-14T09:47:28-0500

While ascending:


"F=m(g+a)=5280\text{ lbs}·\text{ft}/\text s^2."


While descending:


"F=m(g-a)=4960\text{ lbs}·\text{ft}/\text s^2."